# Specialist vision-language models for clinical ophthalmology

### v1

@@ 2.5 RetinaVLM accurately detects imaging biomarkers to make recommendations
We find that RetinaVLM-Specialist outperforms both LLaVA-Med and Med-Flamingo in the detection of seven out of the ten of main biomarkers related to AMD (see Figure 6b). Biomarkers that were more severe, larger, and more numerous were detected with higher accuracy by RetinaVLM-Specialist than less advanced presentations (see Figure 6c). Most of the smaller biomarkers, such as small amounts of intraretinal fluid, drusen and hyperreflective foci, which can be as small as 30 $\mu m$ in size [21], were detected with lower sensitivity. Overall, clinically important hallmarks of late AMD were detected with a very high sensitivity. Large volumes of subretinal and intraretinal fluid were detected in 80% and 78% of cases, respectively, and severe levels of hypertransmission in 84% of cases.
Finally, to visualize the functioning of RetinaVLM-Specialist, we calculated saliency maps based on Grad-CAM [22]. These saliency maps highlight the image regions deemed most important by the model when writing specific passages of the report (see Figure 6d). We refer to Figure 14 for four additional image reports with corresponding saliency maps. We qualitatively found that RetinaVLM-Specialist is influenced by different imaging biomarkers when writing different passages of the report, and in making its final recommendation. We observed the saliency maps were especially effective for highlighting hyperreflective material, RPE irregularities and hypertransmission.
@@ 3 Discussion
In this study, we have presented RetinaVLM, the first specialist generative visual language model in medicine. Given a retinal OCT image, RetinaVLM provides accurate, detailed textual responses related to disease staging, referral or biomarker identification of AMD. While large foundation deep learning models have been employed for retinal image analysis before [23, 18], our generative VLM is the first model that can flexibly process varied textual queries related to complex ophthalmological decisions and return detailed written responses. Through the use of language as primary communication medium, artificial intelligence systems are able to dynamically perform new tasks and meet the evolving requirements of image-based clinical decision makers.
@@
In extensive experiments, RetinaVLM significantly outperformed state-of-the-art generative VLMs designed for medical use. We found that these are unable to interpret OCT images, derive the AMD disease stage and follow standard referral guidelines for AMD. Specifically, we have shown that in disease staging, RetinaVLM surpasses LLaVA-Med, the most performant open-source medical VLM, and is approaching the accuracy of junior ophthalmologists. When testing the ability of VLMs to screen for high-risk patients, LLaVA-Med substantially underperforms compared to junior ophthalmologists and even non-specialist opticians. In comparison, RetinaVLM-Specialist’s reports reduced the number of incorrect urgent referrals by almost four times compared to opticians and had higher recall for urgent referrals than junior ophthalmologists. Finally, RetinaVLM is able to reinforce its decisions by citing observable biomarkers within the written report, and highlighting their corresponding regions within the image.
@@
We postulate that the poor performance of existing medical foundation models stems from their lack of detailed knowledge related retinal OCT and AMD. Current VLMs are trained on broad, unstructured datasets that are extracted from medical textbooks, scientific publications or social media posts of healthcare professional [9]. In the United Kingdom and the United States, clinical trainees aspiring to become specialists must undergo up to ten years of post-graduate training to obtain the grade of a board-certified consultant. We argue that training data of current foundation medical VLMs lacks this specialist knowledge and experience, hindering their effective application to real-world clinical tasks.
@@
A core innovation of our work was the creation of a dedicated training curriculum that specializes VLMs in image-based clinical decision making. Analogously to current medical education, this curriculum deconstructs clinical problems into sets of mandatory capabilities required for their resolution and selectively trains VLMs in these skills. To this end, we obtained a large number of tabular reports by processing of retrospectively collected clinical data using advanced algorithms. Additionally, we tasked ophthalmologists to produce a limited number of highly specific textual reports. In total, our curriculum comprises 41,926 OCT images with 479,710 corresponding visual questions and answers. While still modest in size compared to substantially larger foundation datasets, we believe such curated needs-driven approaches are required to deploy language models specialist healthcare. In a similar vein, leading technology companies in artificial intelligence have also started to look beyond the internet’s image and text data to source specialized training data to train LLMs and VLMs in disciplines such as computer programming, journalism, mathematics [24, 25].
@@
Naturally, the quality of our curriculum depends on the underlying reports, in particular that of the 330 collected detailed textual reports. The majority of the reports used to create RetinaVLM-Specialist were written by a junior ophthalmologist with three years experience. The remaining reports were written by an ophthalmologist with ten years of experience, and their reports were more comprehensive. To match or surpass the performance of the average junior ophthalmologist, we seek to finetune this model further on a small number of reports written by intermediate and senior ophthalmologists.
Additionally, the data may reflect local clinical definitions and workflows. While RetinaVLM-Specialist was trained on reports written by the two UK-based ophthalmologists, it was tested on labels derived from four separate ophthalmologists from different countries and hospitals. In our testing we identified differences in the staging definitions used by the UK-based and Austria-based ophthalmologists, especially with regard to identifying fibrovascular features that differentiate intermediate from inactive late wet AMD. This discrepancy resulted in a number of the misclassifications by RetinaVLM.
@@
We found that the LLM generating the question-answer pairs from the reports was sensitive to the specifics of the generation instruction. Extensive trial and error were required to arrive at several instructions, listed in Section A.2, that resulted in diverse sets of high quality question-answer pairs. We discern that all aspects of dataset creation - deciding on the required capabilities, collecting specialized annotations and converting these to question-answer pairs - should be formalized to systematically compare different approaches and ultimately scale dataset curation in the future.
Beyond the formalization and extension of the creation of the curriculum, there are other potential technical improvements to RetinaVLM. Currently, RetinaVLM processes a single two-dimensional OCT image from one type of OCT scanner. In ophthalmological practice, decisions are made based on three-dimensional images from multiple time points, although many recent studies on the use of foundation models in ophthalmology also analyze two-dimensional images [18]. We mitigated the impact of this discrepancy on our study by tasking ophthalmologists to select the most relevant two-dimensional slice of the imaged volume before proceeding with the referral decision. In the future, a more sophisticated vision encoder, which is able to handle three-dimensional data from diverse OCT imaging devices, could be integrated with RetinaVLM. Similarly, one may opt to incorporate multimodal information, such as health questionnaires, clinical tests or the patient’s medical history, into the decision making process [26]. The fundamental model architecture and training would remain similar, but the level of reasoning required for differential diagnosis across multiple scans would potentially increase.
@@
RetinaVLM also inherits some of the fundamental limitations of language models. LLMs are prone to confidently present false or fabricated information, termed hallucinations, which has been identified as problematic in medical contexts [13, 14]. RetinaVLM occasionally hallucinates observations of retinal fluid and consequently diagnoses more advanced AMD stages than necessary. RetinaVLM’s output was also sensitive to the wording of questions and instructions. While this had little impact on our qualitative analysis, extensive trial and error was necessary to ensure that RetinaVLM responded with one of the provided options in the quantitative analyses.
@@
In this study, we exclusively trained RetinaVLM for the management of AMD from OCT images, ignoring other retinal pathologies, such as diabetic retinopathy or glaucoma, or imaging modalities, such as color fundus photography [27, 28]. While this enabled us to explore the potential to encode advanced clinical levels of specialism into VLMs at depth, it would limit the applicability of the current version of our models for ocular screening. We hypothesize that VLMs could be specialized on a wider range of ophthalmological tasks by incorporating additional VQA datasets and training VLMs on them. This requires costly and time-consuming curation of specialized training datasets by medical experts. However, we believe that the involvement of medical experts is necessary as routine clinical skills and patient management protocols are rarely documented in existing datasets used to train AI models.
@@ Conclusion
Foundation vision-language models (VLMs) have the potential to revolutionize healthcare by automatically interpreting medical images and communicating their findings in detailed written reports. Trained on large datasets containing millions of medical images and textual annotations, foundation language models have stirred considerable interest for their expert-level performance on medical licensing exams and case studies. However, in this work we have shown that foundation medical VLMs substantially underperform human experts on routine clinical tasks.

### v2

@@ 2.5 RetinaVLM accurately detects imaging biomarkers to make recommendations
We find that RetinaVLM-Specialist outperforms both LLaVA-Med and Med-Flamingo in the detection of seven out of the ten of main biomarkers related to AMD (see Figure 6b). In overall performance, RetinaVLM-Specialist performed similarly well as ChatGPT-4o. In general, biomarkers that were more severe, larger, and more numerous were detected with higher accuracy by RetinaVLM-Specialist than less advanced presentations (see Figure 6c). Most of the smaller biomarkers, such as small amounts of intraretinal fluid, drusen and hyperreflective foci, which can be as small as 30 $\mu m$ in size [29], were detected with lower sensitivity. Overall, clinically important hallmarks of late AMD were detected with a very high sensitivity. Large volumes of subretinal and intraretinal fluid were detected in 80% and 78% of cases, respectively, and severe levels of hypertransmission in 84% of cases. We believe that stronger image encoders capable of detecting smaller features are necessary to improve the performance of RetinaVLM-Specialist on the detection of smaller biomarkers.
Finally, we compute saliency maps which highlight regions of the image that were most important to the model in writing specific passages of the report (see Figure 6d). Qualitatively, we find that RetinaVLM-Specialist attends to relevant regions of the image containing fluid, hypertransmission and retinal pigment epithelium (RPE) irregularities in order to compose passages related to biomarkers and disease stage. Rather than performing biomarker segmentation, for which standard deep segmentation models by [30] would be a more appropriate tool, these maps provide some explanation as to the choice of words and report passages made by the model. We calculate these using Grad-CAM [31] as described in Section 5.7.5, and provide four additional examples in Figure 14.
@@ 3 Discussion
In this study we have presented a curriculum-based approach for the specialization of medical vision-language models that directly leverages the knowledge of domain experts. Given a retinal OCT image, the resulting RetinaVLM-Specialist model generates accurate, detailed textual responses related to disease staging, referral or biomarker identification of AMD. While large foundation deep learning models have been employed for retinal image analysis before [32, 22], our generative VLM is the first model that can flexibly process varied textual queries related to complex ophthalmological decisions and return detailed written responses. Through the use of language as primary communication medium, artificial intelligence systems are able to dynamically perform new tasks and meet the evolving requirements of image-based clinical decision makers.
@@
In extensive experiments, RetinaVLM-Specialist significantly outperformed ChatGPT-4o and two generative VLMs, Med-Flamingo and LLaVA-Med, designed for medical use. RetinaVLM-Base and ChatGPT-4o were more capable in classifying conversion to late stages of AMD than existing medical foundation models, Flamingo and LLaVA-Med. However, they also lacked the ability to identify subtle yet important differences between disease stages and patient risk. Specifically, we have shown that in disease staging, RetinaVLM-Specialist outperforms all baselines and is approaching the accuracy of junior ophthalmologists. Similarly, when testing the ability of VLMs to screen for high-risk patients, ChatGPT-4o outperforms non-specialist opticians on aggregate but significantly underperforms compared to RetinaVLM-Specialist and junior ophthalmologists. In comparison, RetinaVLM-Specialist’s reports reduced the number of incorrect urgent referrals by almost four times compared to opticians and had higher recall for urgent referrals than junior ophthalmologists. Finally, RetinaVLM is able to reinforce its decisions by citing observable biomarkers within the written report, and highlighting their corresponding regions within the image.
@@
We postulate that the poor performance of ChatGPT-4o and medical foundation models alike stems from their lack of detailed knowledge related retinal OCT and AMD. Current VLMs including ChatGPT-4o and LLaVA-Med are trained on broad, unstructured datasets that are extracted from medical textbooks, scientific publications or social media posts of healthcare professional [9]. In the United Kingdom and the United States, clinical trainees aspiring to become specialists must undergo up to ten years of post-graduate training to obtain the grade of a board-certified consultant. Classifying AMD requires identifying subtle yet crucial differences between disease stages, particularly in biomarkers such as hypertransmission, drusenoid versus fibrous pigment epithelial detachments, and subretinal hyperreflective material — none of which ChatGPT-4o reported on. While ChatGPT-4o is capable of explaining these differences in language, our analysis indicates it cannot yet make these distinctions in images. This reinforces our insight that the paired image-text training data of current foundation VLMs lacks specialist and experiential knowledge, hindering their effective application to real-world clinical tasks.
@@
A core innovation of our work was the creation of a dedicated training curriculum that specializes VLMs in image-based clinical decision making. Analogously to current medical education, this curriculum deconstructs clinical problems into sets of mandatory capabilities required for their resolution and selectively trains VLMs in these skills. To this end, we obtained a large number of tabular reports by processing of retrospectively collected clinical data using advanced algorithms. Additionally, we tasked ophthalmologists to produce a limited number of highly specific textual reports. In total, our curriculum comprises 41,926 OCT images with 479,710 corresponding visual questions and answers. While still modest in size compared to substantially larger foundation datasets, we believe such curated needs-driven approaches are required to deploy language models specialist healthcare. In a similar vein, leading technology companies in artificial intelligence have also started to look beyond the internet’s image and text data to source specialized training data to train LLMs and VLMs in disciplines such as computer programming, journalism, mathematics [33, 34]. Future work may study specializing generalist models such as ChatGPT-4o by finetuning them on specific, high-quality curricula such as the one introduced in this study.
@@
Naturally, the quality of our curriculum depends on the underlying imaging reports, and in particular those used to create curriculum part 2. The majority of the reports used to create RetinaVLM-Specialist were written by a junior ophthalmologist with three years experience. The remaining reports were written by an ophthalmologist with ten years of experience, and their reports were more comprehensive. While reports written by both were found to be of high quality, future work could evaluate the benefits of a third training curriculum derived by senior ophthalmologists.
By updating instruction sets, VLMs have the potential to be more adaptable than image-only deep learning approaches to differences in clinical practice between countries. However, it is important to note that our curriculum was derived using UK-specific clinical definitions and workflows. As a result, the errors made by RetinaVLM were found to reflect differences in the biomarker and staging definitions used by the UK-based and Austria-based ophthalmologists involved in the testing of the model. In particular, the model was found to be conservative in classifying fibrovascular features that upgrade intermediate AMD to inactive late wet AMD. Thus, to match or surpass the performance of the average junior ophthalmologist, we also aim to establish a consistent cohort of ophthalmologists for training and testing the model.
@@
Another technical limitation of our approach was the sensitivity of the LLM generating the question-answer pairs to the specifics of its instructions. Extensive trial and error were required to arrive at several instructions, listed in Section A.9, that resulted in diverse sets of high quality question-answer pairs. We discern that all aspects of dataset creation - deciding on the required capabilities, collecting specialized annotations and converting these to question-answer pairs - should be formalized to systematically compare different approaches and ultimately scale dataset curation in the future.
@@
RetinaVLM also inherits some of the fundamental limitations of language models. LLMs are prone to confidently present false or fabricated information, termed hallucinations, which has been identified as problematic in medical contexts [13, 14]. This phenomenon has also been observed in VLMs, where the generated text does not relate to any object or feature that is observable in the image [35, 36]. Similarly, we observed that RetinaVLM occasionally hallucinates the presence of retinal fluid and consequently diagnoses more advanced AMD stages than necessary. RetinaVLM’s output was also sensitive to the wording of questions and instructions. While this had little impact on our qualitative analysis, extensive trial and error was necessary to ensure that RetinaVLM responded with one of the provided options in the quantitative analyses.
Beyond the formalization and extension of the curriculum, there are other potential technical improvements to RetinaVLM. Currently, RetinaVLM processes a single two-dimensional OCT image from one type of OCT scanner. In ophthalmological practice, decisions are made based on three-dimensional images from multiple time points, although many recent studies on the use of foundation models in ophthalmology also analyze two-dimensional images [22]. We mitigated the impact of this discrepancy on our study by tasking ophthalmologists to select the most relevant two-dimensional slice of the imaged volume before proceeding with the referral decision. In the future, a more sophisticated vision encoder, which is able to handle three-dimensional data from diverse OCT imaging devices, could be integrated with RetinaVLM. Our results also indicate that stronger image encoders capable of detecting smaller features are necessary to improve the performance of RetinaVLM-Specialist on the detection of smaller biomarkers. Similarly, one may opt to incorporate multimodal information, such as health questionnaires, clinical tests or the patient’s medical history, into the decision making process [37]. The fundamental model architecture and training would remain similar, but the level of reasoning required for differential diagnosis across multiple scans would potentially increase.
@@
Similarly, we exclusively trained RetinaVLM for the management of a single retinal disease, AMD, ignoring other retinal pathologies such as diabetic retinopathy or glaucoma, or imaging modalities, such as color fundus photography [38, 39]. While this enabled us to explore the potential to encode advanced clinical levels of specialism into VLMs at depth, it would severely limit the applicability of the current version of our models for general ocular screening. It also led the model to classify all cases of retinal fluid as wet AMD, where a few cases showed evidence of diabetic retinopathy and other retinal conditions.
In order to ready VLMs for deployment for ocular screening, future work would need to extend RetinaVLM’s curriculum with domain experts from an expanded range of ophthalmological conditions. This requires costly and time-consuming curation of specialized training datasets by medical experts. However, we believe that this is a necessary investment as routine clinical skills and patient management protocols are rarely documented in existing datasets used to train AI models.
@@ Conclusion
Foundation vision-language models (VLMs) have the potential to revolutionize healthcare by automatically interpreting medical images and communicating their findings in detailed written reports. Trained on large datasets containing millions of medical images and textual annotations, foundation language models have stirred considerable interest for their expert-level performance on medical licensing exams and case studies. However, in this work we have shown that foundation medical VLMs such as ChatGPT-4o substantially underperform compared with human experts on routine clinical tasks.